OpenStudy (anonymous):

As incentive for Gretchen to save the $2,500 she recently inherited, her parents tell her that they will pay her 10% simple annual interest on the amount that she saves. How much money will Gretchen have after 10 years if she saves all of the money she inherited?

OpenStudy (anonymous):

I know the formula I=PRT.. but I don't know how to solve it.

OpenStudy (anonymous):

What is the P, the R, and the T? T=time right?

OpenStudy (anonymous):

P is Percent R is rate t is time

OpenStudy (anonymous):

Okay so plug the values into the formula and I'll help you solve it.

OpenStudy (anonymous):

2,500=.10*2,500*10 ? I think.. :O

OpenStudy (anonymous):

No, your trying to get the Incentive or I for 10 years so it would be I=0.10*2500*10

OpenStudy (anonymous):

Close.. :) Now do I just multiply it all?

OpenStudy (anonymous):

yepp, 0.10*2500*10 so try it (:

OpenStudy (anonymous):

wait i think we may have done something wrong...

OpenStudy (anonymous):

this isn't making sense..she should have more but after you multiply it all out I get 2500

OpenStudy (anonymous):

Yeah, don't ask me I mess this up bad. Here are the answer choices: a) $2,708 b) $2,727 c) $2,78 d) $5,000

OpenStudy (anonymous):

c) $2,778*

OpenStudy (anonymous):

this doesn't add up, if I=PRT and P=Percent, R=Rate, and T=Time, it should be set up how we had it I=.10*2500*10 but when you do that you get 2500

OpenStudy (anonymous):

Well, it doesn't matter because you could change it a million times and it still all equal 2,500. IM SO LOST :c

OpenStudy (anonymous):

what does the I stand for?

OpenStudy (anonymous):

Interest?

OpenStudy (anonymous):

Oh okay I get it now, you have I which is the accumulated amount of interest. To find the amount of interest on $2500 over a 10 year period with and annual rate of 10% you use the formula which does give you $2500. the I=$2500 is ONLY the amount of interest, you have to add that to the 2500 she started with. 2500+2500=? and thats your answer

OpenStudy (anonymous):

you get it now?

OpenStudy (anonymous):

yes, thanks:)

OpenStudy (anonymous):

haha no problem(:

OpenStudy (anonymous):

http://oi46.tinypic.com/24buuyx.jpg Help me with another?

OpenStudy (anonymous):

sure, that pic didnt come up tho it said 404 not found

OpenStudy (anonymous):

http://tinypic.com/view.php?pic=24buuyx&s=6

OpenStudy (anonymous):

idk its not coming up on here, can you attach it as a file on here ?

OpenStudy (anonymous):

okay so whats the SA formula of a cube? of a square pyramid?

OpenStudy (anonymous):

I have no idea.

OpenStudy (anonymous):

haha well how many faces does a cube have?

OpenStudy (anonymous):

8..?

OpenStudy (anonymous):

Oh, no a cube has 4 I think.. xD

OpenStudy (anonymous):

plus to one on top and one on bottom so theres 6 in total which makes the SA formula for a cube 6s^2 where s is the length of the side

OpenStudy (anonymous):

It's not a cube!

OpenStudy (anonymous):

do 6(7*6)=SA of the bottom figure

OpenStudy (anonymous):

@srw1496 It's not SA formula!

OpenStudy (anonymous):

what are you talking about? she's trying to find SA and to find the SA of the rectangular prism on the bottom you would do (7*6)(7*6)(7*6)(7*6)(7*6)(7*6)=6(7*6)

OpenStudy (anonymous):

" do 6(7*6)=SA of the bottom figure " NOT SA!!!

OpenStudy (anonymous):

thats how i was taught

OpenStudy (anonymous):

Well, I guess you can learn better by: SA = Area of base + Lateral Area

OpenStudy (anonymous):

I don't know what that is...I was taught depending on the shape, to find the area of one face then multiply it by the amount of sides which is 6 in this case that finds the SA of the base shape, then you would find the SA of the top shape, the triangular pyramid and add them together

OpenStudy (anonymous):

Since the base is square, Area of base: 6² = 36 4 other faces are rectangular: ...

OpenStudy (anonymous):

That's just for the Prism, not for pyramid yet!